OpenStudy (anonymous):

Solve the following equation. Be sure to check the answer in the original equation if you multiply both sides by an expression that contains the variable. (Examples 1-6. If there is no solution, enter NO SOLUTION.) (8/x^2-16)-(1/x-4)=(4/x+4)

OpenStudy (anonymous):

want me to put up my answer? i know its wrong though

OpenStudy (mertsj):

multiply both sides by x^2-16 and get: 8 -(x+4)=4(x-4) or 8 - x - 4 = 4x-16 or 4 - x = 4x-16 20=5x, x = 4 Which causes the denominator to be 0 so there is NO SOLUTION

OpenStudy (anonymous):

yea i got the same things on the numerator and was like wtf? ha